# Matthias Schonlau

### v1

@@ 3.4 Parallel univariate plots
A hammock plot shows an alternating sequence of univariate plots (univariate bars on the axes and bivariate plots (connecting bars between two neighboring axes). If we omit the bivariate plots, parallel univariate plots emerge as an edge case. In Stata, we simply specify space(1). Figure 9 shows an example for the 2020 Tour de France data. Without the connecting bars we are able to display more variables. We see that the winner, Pogacar, was not always in the winning group, but except for stage 16 he was never far behind.
@@
Income <<< 15K 47.2
15K $\leq$ income<<< 30K 26.7
30K $\leq$ income<<< 50K 15
@@
The asthma plot corresponding to Table 1 was constructed with the following command:
@@
    >   labelopt(size(vsmall)) xlab(,labsize(small))
The value for option uni__\__fraction (which affects the relative size of the bars) was chosen as large as possible without the two bars corresponding to 0 and 1 comorbidities touching. (If we move a bar upwards, then the variable would become categorical because the meaning of distance is lost.)
@@ 3.5 Re-ordering variables and categories
The variable order can be rearranged to make the overall plot more understandable or to make a specific point. As a starting point, one may arrange the order to maximize the correlation of neighboring variables. For the auto data, Figure 2 shows a variable ordering starting with mileage (on the left), where the following variable is chosen to maximize the correlation with the current variable. For example, among the remaining variables, weight has the highest correlation with mileage, “car origin” has the highest correlation with weight, and so forth.
@@
Figure 11: Hammock plot of the auto data: variables are ordered to minimize the squared correlation among neighboring variables. Instead we prefer the ordering in Figure 2.
@@
When the variable is continuous or ordered categorical, the order of the values and categories is fixed (in rare circumstances it may be sensible to reverse the order). When a variable is unordered categorical, we can re-arrange the categories to make the graph more readable or to bring out an aspect of the story we wish to tell. Often, a graph is easier to read when — among neighboring variables — low bars connect to low bars and high bars connect to high bars.
@@
The only unordered categorical variable in Figure 2 is car origin (foreign vs domestic). In the current order, connecting bars from domestic (at the bottom) connect to low repair bars (at the bottom). This accentuates the finding that domestic cars have worse repair records.
4 Conclusion
@@
The hammock plot visualizes categorical and numerical variables in the same plot. The plot is often most effective when a category or variable range is highlighted. Among the variations of the hammock plot is the alluvial plot, developed later and unable to accommodate numerical variables. The hammock plot or one of its variations is not the only way to gain insight into data. But it is one more tool in your toolbox and it is often useful.
The parallel univariate plots in Figures 9 and 10 displayed bars. Future work may extend such parallel univariate plots by displaying a stacked bar chart for each x-variable instead of the bars, or a boxplot/ violinplot for continuous variables such as age. Parallel univariate plots that do not use highlighting can be constructed from marginal statistics alone (e.g., Figure 10). Parallel univariate plots that are linked through highlighting such as Figure 9 require bivariate statistics (e.g., variables “rider” and the stage time “s1” for stage 1). Because we usually want to experiment with highlighting more than one variable, parallel univariate plots that use highlighting effectively require the whole data set.
I (Schonlau) have developed the hammock command since 2002. For the most recent version see the Boston SSC archive (ssc install hammock) or my github page (Schonlau, 2025).

### v2

@@ 4.4 Parallel univariate plots
A hammock plot shows an alternating sequence of univariate plots (univariate bars on the axes and bivariate plots (connectors between two neighboring axes). If we omit the bivariate plots, parallel univariate plots emerge as an edge case. In Stata, we simply specify space(1). Figure 8 shows an example for the 2020 Tour de France data. Without the connectors we are able to display more variables. We see that the winner, Pogacar, was not always in the winning group, but except for stage 16 he was never far behind.
@@
Income << 15K 47.2
15K $\leq$ income<< 30K 26.7
30K $\leq$ income<< 50K 15
@@
The plot corresponding to Table 1 was constructed with the following command:
@@
    >   colorlist(orange%20) ///
    >   labelopt(size(vsmall))  ///
    >   xlab(,labsize(small))
The option labelopt(size(vsmall) reduces the font size of the labels on the axes; the option xlab(,labsize(small) reduces the font size of the variable names. The grayscale version of this plot uses the option colorlist(gs12). The value for option uni_\_fraction (which affects the relative size of the univariate bars) was chosen as large as possible without the two univariate bars corresponding to 0 and 1 comorbidities touching. (Spacing for numerical variables is not arbitrary: If one were to move a univariate bar for comorbidities upwards, then the variable would become categorical because the meaning of distance is lost.)
@@ 4.5 Re-ordering variables and categories
The variable order can be rearranged to make the overall plot more understandable or to make a specific point. As a starting point, one may arrange the order to maximize the squared correlation of neighboring variables (or another measure of association for categorical variables). For the auto data, Figure 2 shows a variable ordering starting with mileage (on the left), where the following variable is chosen to maximize the squared correlation with the current variable. For example, among the remaining variables, weight has the highest squared correlation with mileage, “car origin” has the highest squared correlation with weight, and so forth.
@@
Figure 10: Hammock plot of the auto data: variables are ordered to minimize the squared correlation among neighboring variables. Instead, we prefer the ordering in Figure 2.
@@
When the variable is continuous or ordered categorical, the order of the values and categories is fixed (in rare circumstances it may be sensible to reverse the order). When a variable is unordered categorical, we can re-arrange the categories to make the graph more readable or to bring out an aspect of the story we wish to tell. Often, a graph is easier to read when — among neighboring variables — low univariate bars connect to low univariate bars and high univariate bars connect to high univariate bars.
@@
The only unordered categorical variable in Figure 2 is car origin (foreign vs domestic). In the current order, connectors from domestic (at the bottom) connect to low repair univariate bars (at the bottom). This accentuates the finding that domestic cars have worse repair records.
5 Conclusions
The hammock plot visualizes categorical and numerical variables within a single, unified plot. By mapping categories to parallel vertical axes and connecting observations across variables, the hammock plot supports simultaneous inspection of distributions, relationships, and individual trajectories. This is useful for both analysis and communication and makes hammock plots applicable in a wide range of settings.
In exploratory data analysis, hammock plots enable analysts to quickly assess relationships among multiple variables. Without highlighting, univariate and select bivariate relations (those of neighboring variables) are shown. With highlighting, additional bivariate and select tri-variate relationships are also shown. There is a dual focus on distributions and highlighting individual cases; and this distinguishes hammock plots from traditional summary graphics such as bar charts or boxplots.
Hammock plots are effective for the visual identification of outliers. Unlike numerical detection methods, this visual approach allows analysts to judge outliers in the context of multiple variables simultaneously, which is particularly useful when deviations are meaningful only in relation to several variables.
Hammock plots are useful in uncovering missing value patterns. At a glance, analysts see the extent of missingness for each variable. Highlighting missing values of one variable, analysts can determine, for example, whether missingness across variables is mostly due to the same subset of observations. If not, such highlighting may reveal a more intricate missing value pattern such as a dependency on another variable.
Hammock plots are well suited for comparative analysis. Highlighting a group of observations shows whether the observations are systematically different than the remainder across a range of variables. For example, highlighting is useful to compare outcomes for a treatment versus a control group.
Hammock plots support tracing individual observations across multiple variables. This is particularly intuitive for longitudinal data or when the sequence of variable is ordered in time or otherwise meaningful. This capability supports visual story telling and helps connect statistical summaries to concrete cases.
Hammock plots give a visual summary of the data and can serve as an alternative to tables. While tables provide exact values, hammock plots visually distinguish categories of different sizes and can showcase relationships. This makes them especially useful when communicating findings to audiences interested in patterns rather than precise numerical detail.
@@ 5 Conclusions
Among the variations of the hammock plot is the alluvial plot, developed later and unable to accommodate numerical variables. The hammock plot (or one of its variations) is not the only way to gain insight into data. It is, however, one more tool in your toolbox and it is often useful.
In summary, hammock plots are a flexible visualization technique that bridges summary-level insight and observation-level detail, supporting both exploration and communication for multiple mixed categorical/numerical variables.